HUBBLE SPACE TELESCOPE DAILY REPORT #1474 PERIOD COVERED: 0000Z (UTC) 09/12/95 - 0000Z (UTC) 09/13/95 Daily Status Report as of 256/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6274 (Linking the HIPPARCOS to the Extragalactic Reference Frame) The FGS #3 was used to make astrometric observations of extra-galactic object (EGO) 2200+420, and reference stars 1040-REF and INCA221-17 in order to link the HIPPARCOS satellite to the extragalactic reference frame. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrent with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6095 (Hot Stellar Populations in Globular Cluster Cores) The WF/PC-2 was used to make observations of globular cluster NGC-6441 to study the radial gradients of the hot stellar population in its core region. Observations were made in the blue, violet and far ultraviolet. The observations were completed as planned, and no problems were reported. 1.4 Completed FGS 6175 (Three Points of Light Calibration) The FGSs were used in the Trans Mode to make reference transfer function astrometric observations of standard star UPGREN69 to make a three points of light calibration. Observations were made using three different filters at three different locations in the "pickle". The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky survey observations at latitudes larger than 10 degrees taken in parallel with the prime science observations. The observations were executed as scheduled, and no anomalies were noted 1.6 Completed WF/PC-2 5957 (Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make observations for radio galaxy 3C231.0 to study the morphology of the radio galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of GHRS/2 5861 (The Boron Abundances of Tidily Locked Binaries in the Hyades Cluster) The High Resolution Spectrograph (Side Two) was used to make observations of the tidily locked short period binary stars HD28205 and HD28033 which are located in the Hyades star cluster in order to measure their boron abundance. The observations were completed as scheduled, and no anomalies were reported. 1.8 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal Monitor) The WF/PC-2 was used to make internal bias, internal flat images and K-Spot calibrations to monitor the health of the cameras. The observations were executed as scheduled, and no problems were reported. 1.9 Completed FGS 6151 (Masses of Pre-main Sequence Stars: Cycle 4 GO/CAL) The FGSs were used to make observations of the pre-main sequence binary star HV-Tau located in the Taurus star forming region. The FGS data will provide the apparent orbits needed to compute the total mass of the system. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed FGS 6051 (Dynamical Masses For The Stars in the Pre-Main Sequence Spectroscopic Class) The FGS were used to make astrometric measurements of the location of star STAR-04560 (aka, GSC2387+535) to measure its relative orbit in a pre-main sequence binary system. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 One acquisition at 256/091152Z failed due to the search radius limit being exceeded. HSTAR 5285 was written to document this event. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: The STR dump performed at 256/0021Z had five correctable bit slips. A redump, which occurred at 256/0207Z had 14 bit slips. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-